OpenStudy (anonymous):

sqrt2x^2+sqrt3x+1=0 use the discriminant to determin how many real roots each equation

OpenStudy (anonymous):

that is not even right!

OpenStudy (anonymous):

\[b^2-4ac=\sqrt{3}^2-4\times \sqrt{2}\times 1=3-4\sqrt{2}\]

myininaya (myininaya):

\[b^2-4ac=(\sqrt{3})^2-4(\sqrt{2})(1)=3-4\sqrt{2}\]

OpenStudy (anonymous):

ok is this the answer. This is what I I wrote for the equation

OpenStudy (anonymous):

I guess i am confused

myininaya (myininaya):

no real roots the discriminant is <0

OpenStudy (anonymous):

you job is to see if this number is positive, negative or zero. it is negative (easy check) so not real solutions

myininaya (myininaya):

>0=> two real roots =0=> one real root (multiplicity 2) <0=> no real roots

OpenStudy (anonymous):

Thank you